- Astronomers discover an exoplanet (a planet of a star other than the Sun) that has an orbital period of 3.63 Earth years in its circular orbit around its sun, which is a star with a measured mass of 3.65×10^30 kg . Find the radius of the exoplanet's orbit.

- In order better to map the surface features of the Moon, a 353 kg imaging satellite is put into circular orbit around the Moon at an altitude of 137 km. Calculate the satellite's gravitational potential energy, and total orbital energy. The radius and mass of the Moon are 1740 km and 7.36×10^22 kg.

Gravitational Potential Energy - ?

Total Orbital Energy - ?
Answer #1

1.

using kepler's 3rd law of planetary motion

T^2 = 4pi^2 R^3 / ( GM)

(3.63* 365* 24* 3600)^2 = 4* 3.14^2* R^3 / ( 6.67* 10^-11* 3.65* 10^30)

R = 2.007* 10^12 m

======

plz post other problem separately

Comment before rate in case any doubt, will reply for sure.. goodluck